OpenStudy (anonymous):

Evaluate the flux integral FdS where F= (1y,5z,5x) and is the surface of the plane 2x+4y+z=8 in the first octant oriented upward.

OpenStudy (amistre64):

i remember something about F.n

OpenStudy (amistre64):

or to be more vague\[\int\vec F\dot{}\frac{\nabla f}{|\nabla f|}\sqrt{something??}\]

OpenStudy (amistre64):

any of this ring a bell?

OpenStudy (amistre64):

for a refresher, try this: http://tutorial.math.lamar.edu/Classes/CalcIII/SurfIntVectorField.aspx

OpenStudy (anonymous):

Have you tried anything? I can explain it but I'm not going to do it for you.

OpenStudy (anonymous):

is the integral \[\int\limits_{?}^{?} (78y+45x-160) \] but idk the bounds.

OpenStudy (anonymous):

Let: \[\vec{\Phi}(x,y)=(x,y,8-2x-4y)\] Find: \[\frac{\partial \vec{\Phi}}{\partial x} \times \frac{\partial \vec{\Phi}}{\partial y}\] Once you have that you take the dot product with: \[\vec{F}(\vec{\Phi}(x,y))\] From your cross product you can determine if its orientation preserving or reversing. And if the plane is bounded in the first octant then: \[z=0 \implies y=\frac{-x}{2}+2 \implies 0\le y \le \frac{-x}{2}+2 \implies 0 \le x \le 4\] I believe because you're looking at the "shadow" of the plane in the x-y plane and you have something that looks like:  I believe.